OpenStudy (anonymous):

Can someone explain what the following question is asking for? Simplify: (dy / dx) + (d²y / dx²) y = e^3x

OpenStudy (anonymous):

You are given with y = e^3x So: Here firstly find dy/dx then find d^2y/dx^2 Then after finding: just add them..

OpenStudy (anonymous):

9x^4 :)

OpenStudy (anonymous):

Nope, not gettin git...

OpenStudy (anonymous):

Do you know what is the derivative of \(\large e^{3x}\)..??

ganeshie8 (ganeshie8):

can u reply water's q

OpenStudy (anonymous):

Answer: yes. What I'm asking is: what is the question asking me to do?

ganeshie8 (ganeshie8):

see water's first reply. just the q is asking u to find first derivative , second derivative, and add them

OpenStudy (anonymous):

so you're saying (d²y / dx²) is another way of writing "second derivative"?

ganeshie8 (ganeshie8):

yep, another way is y''

OpenStudy (unklerhaukus):

\[y(x)=f\big(g(x)\big)=e^{\alpha x}\] \[f(x)=e^x\qquad g(x)=\alpha x\]\[f'(x)=e^x\qquad g'(x)=\alpha\] \[y'(x)=g'(x)f'\big(g(x)\big)=\alpha e^{\alpha x}\]